# QVGen: Pushing the Limit of Quantized Video Generative Models

### v1

@@ Abstract
Video diffusion models (DMs) have enabled high-quality video synthesis. Yet, their substantial computational and memory demands pose serious challenges to real-world deployment, even on high-end GPUs. As a commonly adopted solution, quantization has achieved notable successes in reducing cost for image DMs, while its direct application to video DMs remains ineffective. In this paper, we present QVGen, a novel quantization-aware training (QAT) framework tailored for high-performance and inference-efficient video DMs under extremely low-bit quantization (i.e., 4-bit or below). We begin with a theoretical analysis demonstrating that reducing the gradient norm is essential to facilitate convergence for QAT. To this end, we introduce auxiliary modules ($\Phi$) to mitigate large quantization errors, leading to significantly enhanced convergence. To eliminate the inference overhead of $\Phi$, we propose a rank-decay strategy that progressively eliminates $\Phi$. Specifically, we repeatedly employ singular value decomposition (SVD) and a proposed rank-based regularization ${\bm{\gamma}}$ to identify and decay low-contributing components. This strategy retains performance while zeroing out inference overhead. Extensive experiments across $4$ state-of-the-art (SOTA) video DMs, with parameter sizes ranging from $1.3$B14similar-toabsent14\,$\sim$ 14∼ 14B, show that QVGen is the first to reach full-precision comparable quality under 4-bit settings. Moreover, it significantly outperforms existing methods. For instance, our 3-bit CogVideoX-2B achieves improvements of $+25.28$ in Dynamic Degree and $+8.43$ in Scene Consistency on VBench.
@@ 1 Introduction
Recently, advancements in artificial intelligence-generated content (AIGC) have led to significant breakthroughs in text [60, 7], image [68, 31], and video synthesis [64, 28]. The development of video generative models, driven by the powerful diffusion transformer (DiT) [50] architecture, has been particularly notable. Leading video diffusion models (DMs), such as closed-source OpenAI Sora [49] and Kling [29], and open-source Wan  [63] and CogVideoX [72], can successfully model motion dynamics, semantic scenes, physical consistency, etc. Despite their impressive performance, these models demand high computational resources and substantial peak memory, especially when generating long videos at a high resolution. For example, Wan 14B [64] requires more than 30 minutes and 50GB of GPU memory to generate a 10-second $720p$ resolution video clip on a single $H100$ GPU.
@@
In light of these problems, model quantization, which maps high-precision (e.g., FP16/BF16) data to low-precision (e.g., INT8/INT4) formats, stands out as a compelling solution. For instance, employing 4-bit models with fast kernel implementation can achieve a significant 3×3$\times$3 × speedup ratio with about 4×4$\times$4 × model size reduction compared with floating-point models on NVIDIA RTX4090 GPUs [32]. However, quantizing video DMs is more challenging than quantizing image DMs, and it has not received adequate attention. As shown in Fig. 1, applying prior high-performing approaches [32, 8, 35, 17] to quantize a video DM into ultra-low bits ($\leq$ 4 bits) is ineffective. In contrast to post-training quantization (PTQ), quantization-aware training (QAT) can obtain superior performance through training quantized weights. Nevertheless, it still leads to severe video quality degradation, as demonstrated by Fig. 1 (a) vs. (c)-(e). This highlights the need for an improved QAT framework to preserve video DMs’ exceptional performance under 4-bit or lower quantization.
Specifically, we first provide a theoretical analysis showing that minimizing the gradient norm $\|{\bm{g}}_{t}\|_{2}$ is the key to improving the convergence of QAT for video DMs. Motivated by this finding, we introduce auxiliary modules $\Phi$ for the quantized video DM to mitigate quantization errors. These modules effectively help narrow the discrepancy between the discrete quantized and full-precision models, leading to stable optimization and largely reduced $\|{\bm{g}}_{t}\|_{2}$. The quantized DM thus achieves better convergence. Our observation also implies that the significant performance drops (Fig. 1) of the existing SOTA QAT method [35] may result from its high $\|{\bm{g}}_{t}\|_{2}$ (Fig. 3).
Moreover, to adopt $\Phi$ for improving QAT while avoiding its substantial inference overhead, we progressively remove $\Phi$ during training. Upon further analysis, we have found that the amount of small singular values in ${\mathbf{W}}_{\Phi}$ (the weight of $\Phi$) increases throughout the training process. This indicates that the quantity of low-contributing components in ${\mathbf{W}}_{\Phi}$, which are related to small singular values [75, 71], grows during QAT. As a result, an increasing number of these components can be removed with minimal impact on training. Leveraging this insight, we introduce a rank-decay strategy to progressively shrink ${\mathbf{W}}_{\Phi}$. To be more specific, singular value decomposition (SVD) is first applied to recognize ${\mathbf{W}}_{\Phi}$’s low-impact components. Then, a rank-based regularization ${\bm{\gamma}}$ is utilized to gradually decay these components to $\varnothing$. Such processes (i.e., decompose and then decay) are repeated until ${\mathbf{W}}_{\Phi}$ is fully eliminated, which also means that $\Phi$ is removed. In terms of results, this strategy incurs minimal performance impact while getting rid of the inference overhead.
To summarize, our contributions are as follows:
  • We introduce a general-purpose QAT paradigm, called QVGen. To our knowledge, this is the first QAT method for video generation and achieves effective 3-bit and 4-bit quantization.
  • To optimize extremely low-bit QAT, we enhance a quantized DM with auxiliary modules ($\Phi$) to reduce the gradient norm. Our theoretical and empirical analysis validates the effectiveness of this method in improving convergence.
  • To eliminate the significant inference overhead introduced by $\Phi$, we propose a rank-decay strategy that progressively shrinks $\Phi$. It iteratively performs SVD and applies a rank-based regularization ${\bm{\gamma}}$ to obtain and decay low-impact components of $\Phi$, respectively. As a result, this method incurs minimal impact on performance.
  • Extensive experiments across advancing CogVideoX [72] and Wan [64] families demonstrate the SOTA performance of QVGen. Notably, our W4A4 model is the first time to show full-precision comparable performance. In addition, we apply QVGen to Wan 14B, one of the largest SOTA open-source models, and observe negligible performance drops on VBench-2.0 [78].
2 Related Work
Video diffusion models. Building upon the remarkable success of diffusion models (DMs) [19, 56, 5] in image generation [31, 68], the exploration in the field of video generation [72, 64, 28] is also becoming popular. In contrast to convolution-based diffusion models [22, 4], the success of OpenAI Sora [49] has spurred researchers to adopt the diffusion transformer (DiT) [50] architecture and scale it up for high-quality video generation. However, advanced video DiTs [64, 72, 28] often involve billions of parameters, lengthy multi-step denoising, and intensive computation over long frame sequences. This results in substantial time and memory overhead, which limits their practical deployment. To enable faster video generation, some works have introduced step-distillation [73, 39] on pre-trained models to shorten the denoising trajectory. Others focus on efficient attention [74, 67], feature caching [45, 83], or parallel inference [11, 10] to accelerate per-step computations. Moreover, to achieve memory-efficient inference, existing research has explored efficient architecture design [66, 40], structure pruning [2, 77], and model quantization [76, 6, 59]. These methods aim to achieve both model size and computational cost reduction.
Model quantization. Quantization [26] is a predominant technique for minimizing storage and accelerating inference. It can be categorized into post-training quantization (PTQ) [37, 46] and quantization-aware training (QAT) [61, 82, 8, 12]. PTQ compresses models without re-training, making it fast and data-efficient. Nevertheless, it may result in suboptimal performance, especially under ultra-low bit-width (e.g., 3/4-bit). Conversely, QAT applies quantization during training or finetuning and typically achieves higher compression rates with less performance degradation. For DMs, previous quantization research [35, 32, 17, 34, 18, 23, 24, 55, 65, 54, 63] primarily focuses on image generation. Video DMs, which incorporate complex temporal and spatial modeling, are still challenging for low-bit quantization. QVD [59] and Q-DiT [6] first apply PTQ to convolution-based [70, 13] and DiT-based [81] video DMs, respectively. Furthermore, ViDiT-[76] employs mixed-precision and fine-grained PTQ to improve performance. However, it experiences video quality loss with 8-bit activation quantization and has not been extended to more advanced models [72, 64]. Consequently, QAT for advanced video DMs is urgently needed.
In this work, we identify and address the ineffective low-bit QAT for the video DM. Our proposed method significantly enhances model performance and incurs zero inference overhead in ultra-low-bit settings. Moreover, our framework is orthogonal to existing QAT methods, which target gradient estimation [12, 8], oscillation reduction [47, 35], mixed-precision schemes [61, 51], etc.
@@ 3 Preliminaries
Video diffusion modeling. The video DM [22, 81] extends image diffusion frameworks [35, 56] into the temporal domain by learning dynamic inter-frame dependencies. Let ${\bm{x}}_{0}\in\mathbb{R}^{f\times h\times w\times c}$ be a latent video variable, where $f$ denotes the count of video frames, each of size $h\times w$ with $c$ channels. DMs are trained to denoise samples generated by adding random Gaussian noise ${\bm{\epsilon}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ to ${\bm{x}}_{0}$:
@@
[…]
@@
where $\mathcal{C}$ represents conditional guidance, like texts or images, and $\|\cdot\|_{F}$ denotes the Frobenius norm. Additionally, v-prediction [53] (i.e., predict $\frac{d{\bm{x}}_{\tau}}{d\tau}$) is also a prevailing option [72, 64, 28] as the target. During inference, we can employ ${\bm{\epsilon}}_{\theta}$ with various sampling methods [43, 79] progressively denoising from a random Gaussian noise ${\bm{x}}_{N}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ to a clean video variable. The raw video is obtained by decoding the variable via a video variational auto-encoder (VAE) [72, 14, 64].
Quantization. The current video DM based on the diffusion transformer (DiT) [50] architecture primarily consists of linear layers. Given an input ${\mathbf{X}}\in\mathbb{R}^{m\times k}$, a full-precision linear layer with weight ${\mathbf{W}}\in\mathbb{R}^{n\times m}$ and the layer’s quantized version can be formulated as:
@@
$\begin{array}[]{ll}{\mathcal{Q}}_{b}({\mathbf{X}})=(\mathrm{clamp}(\lfloor\frac{{\mathbf{X}}}{s}\rceil+z,0,2^{b}-1)-z)\times s,\mathrm{where}~{}s=\frac{\max({\mathbf{X}})-\min({\mathbf{X}})}{2^{b}-1},z=-\lfloor\frac{\min({\mathbf{X}})}{s}\rceil.\end{array}$ (4)
[…]
@@
Similar to existing works [35, 80], we employ the full-precision model as the teacher to guide the training of the quantized model in a knowledge distillation-based (KD-based) manner. Therefore, the training loss can be defined as:
$\mathcal{L}=\mathbb{E}_{{\bm{x}}_{0},\mathcal{C},\tau}[\|\hat{\bm{\epsilon}}_{\theta}({\bm{x}}_{\tau},\mathcal{C},\tau)-{\bm{\epsilon}}_{\theta}({\bm{x}}_{\tau},\mathcal{C},\tau)\|_{F}^{2}],$ (6)
@@ 4 QVGen
Figure 2: Overview of the proposed QVGen. (a) This framework integrates auxiliary modules $\Phi$ to improve training convergence (Sec. 4.1). (b) To maintain performance while eliminating inference overhead induced by $\Phi$, we design a rank-decay schedule that progressively shrinks the entire $\Phi$ to $\varnothing$ through iteratively applying the following two strategies (Sec. 4.2): 1) SVD to identify the low-impact components in $\Phi$; 2) A rank-based regularization ${\bm{\gamma}}$ to decay the identified components to $\varnothing$.
Considering the substantial video-quality drops (see Fig. 1) observed in existing QAT methods [35, 17, 8], we believe that the quantized video DM suffers from poor convergence. In the following subsections, we propose QVGen (see Fig. 2) to address this issue while maintaining inference efficiency.
@@ 4.1 Improving Convergence with Auxiliary Modules $\Phi$
To begin with, we analyze the convergence of a quantized video DM using the regret, which is widely used in the analysis of deep learning optimizers [27, 44]. It is defined as:
@@
where $T$ signifies the total number of training iterations and $f_{t}(\cdot)$ is the unknown cost function at iteration $t$. Here, ${\bm{\theta}}_{t}$ represents the parameters of the quantized video DM at training step $t$, constrained within a convex compact set $\mathbb{S}^{d}$, while ${\bm{\theta}}^{*}=\arg\,\min_{{\bm{\theta}}\in\mathbb{S}^{d}}\sum^{T}_{t=1}f_{t}({\bm{\theta}})$ is the optimal parameters. In QAT, ${\bm{\theta}}_{t}$ is updated by gradient descent, with the learning rate $\eta_{t}$ and gradient ${\bm{g}}_{t}$, as:
$\begin{array}[]{ll}{\bm{\theta}}_{t+1}={\bm{\theta}}_{t}-\eta_{t}{\bm{g}}_{t}.\end{array}$ (8)
@@
Assume that $f_{t}$ is convex [^3] This may not hold for deep networks. However, it facilitates the analysis of deep learning models [27, 44]. and $\forall{\bm{\theta}}_{i},{\bm{\theta}}_{j}\in\mathbb{S},\|{\bm{\theta}}_{i}-{\bm{\theta}}_{j}\|_{\infty}\leq D_{\infty}$. Then the average regret is upper-bounded as:
$\begin{array}[]{ll}\frac{R(T)}{T}\leq\frac{dD_{\infty}}{2T\eta_{T}}+\frac{1}{T}\sum^{T}_{t=1}\frac{\eta_{t}}{2}\|{\bm{g}}_{t}\|_{2}^{2}.\end{array}$ (9)
A smaller value of $\frac{R(T)}{T}$ implies a closer convergence to the optimum. Thm. 4.1 (with a proof provided in the appendix) suggests that for a large $T$ (i.e., $\frac{dD_{\infty}}{2T\eta_{T}}$ becomes negligible), minimizing $\|{\bm{g}}_{t}\|_{2}$ is critical for improving convergence behavior of QAT.
A lower $\|{\bm{g}}_{t}\|_{2}$ is typically observed in more stable training processes [58, 69]. Therefore, to reduce $\|{\bm{g}}_{t}\|_{2}$, we aim to stabilize the QAT process by mitigating aggressive training losses (e.g., loss spikes) [30, 33]. Specifically, we introduce a learnable auxiliary module $\Phi$ to enhance each quantized linear layer of a video DM. This module is trained to mitigate severe quantization-induced errors during QAT, thereby preventing aggressive training losses. The forward computation of such a $\Phi$-equipped layer becomes:
@@
(a) CogVideoX-2B [72]
(b) Wan 1.3B [64]
Figure 3: (Upper) $\|{\bm{g}}_{t}\|_{2}$ vs. #steps and (Lower) training loss vs. #steps across different video DMs and 4-bit QAT methods under the same training settings.“$\Phi$” denotes our approach in Sec. 4.1.
To validate the effectiveness of $\Phi$, we conduct experiments for CogVideoX 2B [72] and Wan 1.3B [64]. Compared with the previous SOTA QAT method Q-DM [35], the proposed approach exhibits consistently lower $\|{\bm{g}}_{t}\|_{2}$ and reduced training loss, as depicted in Fig. 3. This aligns well with both the theoretical and the empirical analysis discussed earlier. Therefore, incorporating $\Phi$ in QAT effectively reduces the gradient norm and leads to better convergence for QAT. In addition, as evidenced by Fig. 3, the substantial performance degradation of Q-DM (e.g., depicted in Fig. 1) for the video generation task could be attributed to its relatively large $\|{\bm{g}}_{t}\|_{2}$.

### v3

@@ Abstract
Video diffusion models (DMs) have enabled high-quality video synthesis. Yet, their substantial computational and memory demands pose serious challenges to real-world deployment, even on high-end GPUs. As a commonly adopted solution, quantization has achieved notable successes in reducing cost for image DMs, while its direct application to video DMs remains ineffective. In this paper, we present QVGen, a novel quantization-aware training (QAT) framework tailored for high-performance and inference-efficient video DMs under extremely low-bit quantization (i.e., $4$-bit or below). We begin with a theoretical analysis demonstrating that reducing the gradient norm is essential to facilitate convergence for QAT. To this end, we introduce auxiliary modules ($\Phi$) to mitigate large quantization errors, leading to significantly enhanced convergence. To eliminate the inference overhead of $\Phi$, we propose a rank-decay strategy that progressively eliminates $\Phi$. Specifically, we repeatedly employ singular value decomposition (SVD) and a proposed rank-based regularization $\bm{\gamma}$} to identify and decay low-contributing components. This strategy retains performance while zeroing out inference overhead. Extensive experiments across $4$ state-of-the-art (SOTA) video DMs, with parameter sizes ranging from $1.3\text{B}\sim 14\text{B}$, show that QVGen is the first to reach full-precision comparable quality under $4$-bit settings. Moreover, it significantly outperforms existing methods. For instance, our $3$-bit CogVideoX-22B achieves improvements of +25.28+25.28 in Dynamic Degree and +8.43+8.43 in Scene Consistency on VBench.
@@ 1 Introduction
Recently, advancements in artificial intelligence-generated content (AIGC) have led to significant breakthroughs in text (Touvron et al., 2023; DeepSeek-AI et al., 2025), image (Xie et al., 2025; Labs, 2024), and video synthesis (WanTeam et al., 2025; Kong et al., 2025). The development of video generative models, driven by the powerful diffusion transformer (DiT) (Peebles & Xie, 2023) architecture, has been particularly notable. Leading video diffusion models (DMs), such as closed-source OpenAI Sora (OpenAI, 2024) and Kling (Kuaishou, 2024), and open-source Wan  (WanTeam et al., 2025) and CogVideoX (Yang et al., 2025), can successfully model motion dynamics, semantic scenes, etc. Despite their impressive performance, these models demand high computational resources and substantial peak memory, especially when generating long videos at a high resolution. For example, Wan 1414B requires more than $30$ minutes and 5050GB of GPU memory to generate a $10$-second $720p$ resolution video clip on a single $H100$ GPU.
@@
Figure 1: Comparison of samples generated by CogVideoX-22(Yang et al., 2025) with a fixed random seed. “WxxAyy” denotes “xx”-bit per-channel weight and “yy”-bit per-token activation quantization. Our approach far outperforms previous PTQ (i.e., (f)) and QAT (i.e., (c)-(e)) methods. To be noted, methods (c)-(f) have achieved noticeable performance for $4$-bit image DMs. More visual results can be found in Sec. L.
In light of these problems, model quantization, which maps high-precision (e.g., FP16/BF16) data to low-precision (e.g., INT8/INT4) formats, stands out as a compelling solution. For instance, employing $4$-bit models with fast kernel implementation can achieve a significant $3\times$ speedup ratio with about $4\times$ model size reduction compared with floating-point models on NVIDIA RTX40904090 GPUs (Li et al., 2025). However, quantizing video DMs is more challenging than quantizing image DMs, and it has not received adequate attention. As shown in Fig. 1, applying prior high-performing approaches to quantize a video DM into ultra-low bits ($\leq$ 44 bits) is ineffective. In contrast to post-training quantization (PTQ), quantization-aware training (QAT) can obtain superior performance through training quantized weights. Nevertheless, it still leads to severe video quality degradation, as demonstrated by Fig. 1 (a) vs. (c)-(e). This highlights the need for an improved QAT framework to preserve video DMs’ exceptional performance under $4$-bit or lower quantization.
Specifically, we first provide a theoretical analysis showing that minimizing the gradient norm $\|{\bm{g}}_{t}\|_{2}$ is the key to improving the convergence of QAT for video DMs. Motivated by this finding, we introduce auxiliary modules $\Phi$ for the quantized video DM to mitigate quantization errors. These modules effectively help narrow the discrepancy between the discrete quantized and full-precision models, leading to stable optimization and largely reduced $\|{\bm{g}}_{t}\|_{2}$. The quantized DM thus achieves better convergence. Our observation also implies that the significant performance drops (Fig. 1) of the existing SOTA QAT method (Li et al., 2023b) may result from its high $\|{\bm{g}}_{t}\|_{2}$ (Fig. 3).
Moreover, to adopt $\Phi$ for improving QAT while avoiding its substantial inference overhead, we progressively remove $\Phi$ during training. Upon further analysis, we have found that the amount of small singular values in ${\mathbf{W}}_{\Phi}$ (the weight of $\Phi$) increases throughout the training process. This indicates that the quantity of low-contributing components in ${\mathbf{W}}_{\Phi}$, which are related to small singular values (Zhang et al., 2015; Yang et al., 2020), grows during QAT. As a result, an increasing number of these components can be removed with minimal impact on training. Leveraging this insight, we introduce a rank-decay strategy to progressively shrink ${\mathbf{W}}_{\Phi}$. To be more specific, singular value decomposition (SVD) is first applied to recognize ${\mathbf{W}}_{\Phi}$’s low-impact components. Then, a rank-based regularization $\bm{\gamma}$} is utilized to gradually decay these components to $\varnothing$. Such processes (i.e., decompose and then decay) are repeated until ${\mathbf{W}}_{\Phi}$ is fully eliminated, which also means that $\Phi$ is removed. In terms of results, this strategy incurs minimal performance impact while getting rid of the extra inference overhead.
To summarize, our contributions are as follows:
@@
[…]
@@ 2 Preliminaries
Video diffusion modeling. The video DM (Ho et al., 2022; Zheng et al., 2024) extends image diffusion frameworks (Li et al., 2023b; Song et al., 2021a) into the temporal domain by learning dynamic inter-frame dependencies. Let ${\bm{x}}_{0}\in\mathbb{R}^{f\times h\times w\times c}$ be a latent video variable, where $f$ denotes the count of video frames, each of size $h\times w$ with $c$ channels. DMs are trained to denoise samples generated by adding random Gaussian noise ${\bm{\epsilon}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ to ${\bm{x}}_{0}$:
@@
where $\alpha_{\tau},\sigma_{\tau}>0$ are scalar values that collectively control the signal-to-noise ratio (SNR) according to a given noise schedule (Song et al., 2021b) at timestep $\tau\in[1,\ldots,N]$ [^1] NN denotes the maximum timestep.. One typical training objective (i.e., predict the noise (Ho et al., 2020b)) of a denoiser ${\bm{\epsilon}}_{\theta}$ with parameter $\theta$ can be formulated as follows:
@@
where $\mathcal{C}$ represents conditional guidance, like texts or images, and $\|\cdot\|_{F}$ denotes the Frobenius norm. Additionally, v-prediction (Salimans & Ho, 2022) (i.e., predict $\frac{d{\bm{x}}_{\tau}}{d\tau}$) is also a prevailing option (Yang et al., 2025; WanTeam et al., 2025; Kong et al., 2025) as the target. During inference, we can employ ${\bm{\epsilon}}_{\theta}$ with various sampling methods (Lu et al., 2022; Zheng et al., 2023) progressively denoising from a random Gaussian noise ${\bm{x}}_{N}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ to a clean video variable. The raw video is obtained by decoding the variable via a video variational auto-encoder (VAE) (Yang et al., 2025).
Quantization. The current video DM based on the diffusion transformer (DiT) (Peebles & Xie, 2023) architecture primarily consists of linear layers. Given an input ${\mathbf{X}}\in\mathbb{R}^{m\times k}$, a full-precision linear layer with weight ${\mathbf{W}}\in\mathbb{R}^{n\times m}$ and the layer’s quantized version can be formulated as:
@@
$\begin{array}[]{ll}{\mathcal{Q}}_{b}({\mathbf{X}})=(\mathrm{clip}(\lfloor\frac{{\mathbf{X}}}{s}\rceil+z,0,2^{b}-1)-z)\times s,\mathrm{where}~s=\frac{\max({\mathbf{X}})-\min({\mathbf{X}})}{2^{b}-1},z=-\lfloor\frac{\min({\mathbf{X}})}{s}\rceil.\end{array}$ (4)
Here, quantization parameters $s$ and $z$ denote the scaler and zero shift, respectively. $\mathrm{clip}(\cdot,\cdot,\cdot)$ bounds the integer values into $0,2^{b}-1]$. To ensure the differentiability of the rounding function $\lfloor\cdot\rceil$ for QAT, straight-through estimator (STE) (Bengio et al., 2013) is widely applied as:
@@
Similar to existing works (Li et al., 2023b; Zheng et al., 2025b), we employ the full-precision model as the teacher to guide the training of the quantized model in a knowledge distillation-based (KD-based) manner. Therefore, the training loss can be defined as:
$\mathcal{L}=\mathbb{E}_{{\bm{x}}_{0},\mathcal{C},\tau}[\|\hat{\bm{\epsilon}}_{\theta}({\bm{x}}_{\tau},\mathcal{C},\tau)-{\bm{\epsilon}}_{\theta}({\bm{x}}_{\tau},\mathcal{C},\tau)\|_{F}^{2}].$ (6)
@@ 3 QVGen
Considering the substantial video-quality drops (see Fig. 1) observed in existing QAT methods (Li et al., 2023b; He et al., 2024; Esser et al., 2020b), we believe that the quantized video DM suffers from poor convergence. In the following subsections, we propose QVGen (see Fig. 2) to address this issue while maintaining inference efficiency.
@@ 3.1 Improving Convergence with Auxiliary Modules $\Phi$
Figure 2: Overview of the proposed QVGen. (a) This framework integrates auxiliary modules $\Phi$ to improve training convergence (Sec. 3.1). (b) To maintain performance while eliminating inference overhead induced by $\Phi$, we design a rank-decay schedule that progressively shrinks the entire $\Phi$ to $\varnothing$ through iteratively applying the following two strategies (Sec. 3.2): (i) SVD to identify the low-impact components in $\Phi$; (ii) A rank-based regularization $\bm{\gamma}$} to decay the identified components to $\varnothing$. Overall detailed algorithm can be found in Sec. B
To begin with, we analyze the convergence of a quantized video DM using the regret, which is widely used in analyses of deep learning optimizers (Kingma & Ba, 2017; Luo et al., 2019). It is defined as:
@@
where $T$ signifies the total number of training iterations and $f_{t}(\cdot)$ is the unknown cost function at iteration $t$. Here, ${\bm{\theta}}_{t}$ represents the parameters of the quantized video DM at training step $t$, constrained within a convex compact set $\mathbb{S}^{d}$, while ${\bm{\theta}}^{*}=\arg\,\min_{{\bm{\theta}}\in\mathbb{S}^{d}}\sum^{T}_{t=1}f_{t}({\bm{\theta}})$ is the optimal parameters. In QAT, ${\bm{\theta}}_{t}$ is updated by gradient descent, with the learning rate ${\bm{\eta}}_{t}$ and gradient ${\bm{g}}_{t}$, as:
$\begin{array}[]{ll}{\bm{\theta}}_{t+1}={\bm{\theta}}_{t}-{\bm{\eta}}_{t}{\bm{g}}_{t}.\end{array}$ (8)
@@
[…]
A smaller value of $\frac{R(T)}{T}$ implies a closer convergence to the optimum. Thm. 3.1 (with a proof provided in Sec. C) suggests that for a large $T$ (i.e., $\frac{dD_{\infty}^{2}}{2T\eta_{T}^{m}}$ becomes negligible), minimizing $\|{\bm{g}}_{t}\|_{2}$ is critical for improving convergence behavior of QAT.
A lower $\|{\bm{g}}_{t}\|_{2}$ is typically observed in more stable training processes (Takase et al., 2024; Xie et al., 2024). Therefore, to reduce $\|{\bm{g}}_{t}\|_{2}$, we aim to stabilize the QAT process by mitigating aggressive training losses (e.g., loss spikes) (Kumar et al., 2025; Li et al., 2024b). Specifically, we introduce a learnable auxiliary module $\Phi$ to enhance each quantized linear layer of a video DM. This module is trained to mitigate severe quantization-induced errors during QAT, thereby preventing aggressive training losses. The forward computation of such a $\Phi$-equipped layer becomes:
@@
(a) CogVideoX-22(Yang et al., 2025)
(b) Wan 1.31.3(WanTeam et al., 2025)
[…]
To validate the effectiveness of $\Phi$, we conduct experiments for CogVideoX 22(Yang et al., 2025) and Wan 1.31.3(WanTeam et al., 2025). Compared with the previous SOTA QAT method Q-DM (Li et al., 2023b), the proposed approach exhibits consistently lower $\|{\bm{g}}_{t}\|_{2}$ and reduced training loss, as depicted in Fig. 3. This aligns well with both the theoretical and the empirical analysis discussed earlier. Therefore, incorporating $\Phi$ in QAT effectively reduces the gradient norm and leads to better convergence for QAT. In addition, as evidenced by Fig. 3, the substantial performance degradation of Q-DM (e.g., depicted in Fig. 1) for the video generation task could be attributed to its relatively large $\|{\bm{g}}_{t}\|_{2}$. Besides, we provide further analyses of $\|{\bm{g}}_{t}\|_{2}$ in video generation QAT in Sec. G.